given:
triangle ABC
Perimeter = 15
height = x+1
base = x
find the value of x
please help me answer this problem:(

Answer by ReadingBoosters(3246)

p = 15 = (x+1) + x + third side
Without the third side, perimeter cannot be determined by base and height alone.
If area = 15 then,
a = 15 = 1/2(x)(x+1)
1/2x^2+1/2x = 15
.5x^2 + .5x - 15 = 0
(.5x+3)(x-5)
x = -3/.5 NO
x = 5